As a manager, you may be tasked with making recommendations as to how your organization should structure its transfer pricing. This is especially true in cases when both variable or full cost transfer pricing is acceptable, and the choice is not obvious.

GenTrac Global Inc. manufactures robotic controllers in Division A, a country with a 30% income tax rate, and transfers them to Division B, a country with a 40% income tax. An import duty of 15% of the transfer price is paid on all imported products. The import duty is not deductible in computing taxable income. The controller’s full cost is $1,800 and variable cost is $1,000; they are sold by Division B for $2,100. The tax authorities in both countries allow firms to use either variable cost or full cost as the transfer price.
Answer #1

If Variable cost of transfer pricing is used:

For Division A

Profit = $1000 - $1800 = ($800)

Tax saving on loss = ($800) * 30% = ($240)

Post tax loss = ($800) - ($240) = ($560)

For Division B

Import duty = 15% * $1000 = $150

Profit of Division B ignoring import duty = $2100 - $1000 = $1100

Tax expense = 40% of $1100 = $440

Therefore Post tax profit = $1100 - $150 - $440 = $510

Net profit of the GenTrac Global Inc. = ($560) + $510 = ($50)

If Full cost of transfer pricing is used:

For Division A

Profit = $1800 - $1800 = $0

For Division B

Import duty = 15% * $1800 = $270

Profit of Division B ignoring import duty = $2100 - $1800 = $300

Tax expense = 40% of $300 = $120

Therefore Post tax profit = $300 - $270 - $120 = ($90)

Net profit of the GenTrac Global Inc. = 0 + ($90) = ($90)

Conclusion

As it can be clearly seen that both the options result in loss for the corp, the project is not viable. However when transfer pricing is kept at variable cost then the extent of loss is lower.